A monopolist firm has the following total cost function and demand function and show your calculation.

Quantity demand price total cost

5 8 20

6 7 22

7 6 25

8 5 29

9 4 34

10 3 40

Explain what price will be charged and what output will be produced.   



1
Expert's answer
2021-08-16T15:31:07-0400



"TR=P\times Q"


"MR=\frac{\Delta TR}{\Delta Q}"


"MC=\frac{\Delta TC}{\Delta Q}"


Profit is maximized for monopolist when MR=MC


Condition is satisfied from the table at Q=6 and P=7

MR=MC=2


Therefore at this price and quantity, the profit for monopolist firm is maximised.


therefore

monopolist will charge price P=7

and produce output, Q=6